This is a triganometry problem using Vectors length and direcion. In a regatta, a sailboat sailed 33 km west and 73 km south from the port, and then took the shortest route back. Find the distance and southern most angle in the triangle.
I did find the distance which is 80.1 km.

Answer by Alan3354(69443)

This is a triganometry problem using Vectors length and direcion. In a regatta, a sailboat sailed 33 km west and 73 km south from the port, and then took the shortest route back. Find the distance and southern most angle in the triangle.
I did find the distance which is 80.1 km.
-------------------
It's a right triangle, so use Pyth to find the hypotenuse.
d^2 = 33^2 + 73^2
d^2 = 1089 + 5329 = 6418
d = ~ 80.11 km
----------------
The tangent of the southern angle is 33/73
The angle is ~ 24.33º